OpenStudy (anonymous):

Can someone please help me with Calculus 1 Maximum Volume. Determine the dimensions of a rectangular solid (with a square base) with maximum volume if its surface area is 337.5 square centimeters

OpenStudy (anonymous):

whats the equation for the surface area of a rectangular solid with a square base?

OpenStudy (anonymous):

if u draw it that might help

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

alright lets say the length and width of the base of the box equals x and the height equals y. Since the surface area is the area of all the faces added up then the surface area of this rectangle would b 2x^2 for the base and the top 4xy for the four faces going around so 4xy and 2x^2=Surface area and u know surface area equals 337.5 so 4xy+2x^2=337.5

OpenStudy (anonymous):

how do you know length and width is the same?

OpenStudy (anonymous):

because the base is a square

OpenStudy (anonymous):

oh right thanx

OpenStudy (anonymous):

now u kno that volume equals length*width*height so for our rectangle itll b V=yx^2

OpenStudy (anonymous):

to find the maximum volume u want to differentiate the volume equation but u cant because there r 3 variables. so u need to use ur Surface area equation solve for a variable and then replace that variable in ur volume equation

OpenStudy (anonymous):

so solve for y in the surface area equation i gave u